OpenStudy (anonymous):

Find the equation of a straight line that passes through: (0, 4) and (3,10)

OpenStudy (anonymous):

Using y - y1 x - x1 ----- = ---- y2 - y1 x2 - x1

OpenStudy (anonymous):

do u mean y2-y1 m=------ x2-x1 ?

OpenStudy (anonymous):

y=2x+4 is ur answer

OpenStudy (anonymous):

where did you get the 4 from, can you explain how to you got it?. I appreciate it!

OpenStudy (anonymous):

slope intercept form of a line: y=mx+b i assume u kno how i got m=2 so then our equation is y=2x+b now we need to solve for b so we use 1 of our points and input the values. im gonna use (0,4) cuz its the easiest 4=2(0)+b 4=b so y=2x+4

OpenStudy (anonymous):

Many Thanks, so using your explanation can you tell me if this is right? for the next question (2, 1) and (-2, -7) y = -8x + b 1 = -8(1) + b 1 = -8 +b 8 + 1 = b 9 = b y = -8x + 9

OpenStudy (anonymous):

isshan answered that on ur post